# REBEL: Reinforcement Learning via Regressing Relative Rewards

### v1

@@ Abstract
While originally developed for continuous control problems, Proximal Policy Optimization (PPO) has emerged as the work-horse of a variety of reinforcement learning (RL) applications including the fine-tuning of generative models. Unfortunately, PPO requires multiple heuristics to enable stable convergence (e.g. value networks, clipping) and is notorious for its sensitivity to the precise implementation of these components. In response, we take a step back and ask what a minimalist RL algorithm for the era of generative models would look like. We propose REBEL, an algorithm that cleanly reduces the problem of policy optimization to regressing the relative rewards via a direct policy parameterization between two completions to a prompt, enabling strikingly lightweight implementation. In theory, we prove that fundamental RL algorithms like Natural Policy Gradient can be seen as variants of REBEL, which allows us to match the strongest known theoretical guarantees in terms of convergence and sample complexity in the RL literature. REBEL can also cleanly incorporate offline data and handle the intransitive preferences we frequently see in practice. Empirically, we find that REBEL provides a unified approach to language modeling and image generation with stronger or similar performance as PPO and DPO, all while being simpler to implement and more computationally tractable than PPO.
@@ 1 Introduction
The generality of the reinforcement learning (RL) paradigm is striking: from continuous control problems (Kalashnikov et al., 2018) to, recently, the fine-tuning of generative models (Stiennon et al., 2022; Ouyang et al., 2022), RL has enabled concrete progress across a variety of decision-making tasks. Specifically, when it comes to fine-tuning generative models, Proximal Policy Optimization (PPO, Schulman et al. (2017)) has emerged as the de-facto RL algorithm of choice, from language models (LLMs) (Ziegler et al., 2020; Stiennon et al., 2022; Ouyang et al., 2022; Touvron et al., 2023) to image generative models (Black et al., 2023; Fan et al., 2024; Oertell et al., 2024).
@@
If we take a step back however, it is odd that we are using an algorithm designed for optimizing two-layer networks for continuous control tasks from scratch for fine-tuning the billions of parameters of modern-day generative models. In the continuous control setting, the randomly initialized neural networks and the possible stochasticity in the dynamics necessitate variance reduction through a learned value function as a baseline (Schulman et al., 2015b), while clipping updates is important to limit distribution shift from iteration to iteration (Kakade and Langford, 2002). This means that when applied to generative model fine-tuning, we need to store four models in memory simultaneously (the policy, the reference policy, the critic, and the reward model), each with billions of parameters. Furthermore, we often add a KL regularization to the base model for fine-tuning, making explicit clipping unnecessary nor advisable, as pointed out by Ahmadian et al. (2024). Even outside of the generative modeling context, PPO is notorious for the wide range of performances measured, with differences being attributed to seemingly inconsequential implementation details (Henderson et al., 2019; Engstrom et al., 2020). This begs the question:
@@
Figure 1: We present REBEL: a simple and scalable RL algorithm that performs policy optimization via iteratively regressing the difference in rewards directly in terms of the policy. This allows us to eliminate much of the complexity (e.g. value functions, clipping) of algorithms like PPO (Schulman et al., 2017). We apply REBEL to problems in both image generation and language modeling and find that despite its conceptual and implementation-level simplicity, REBEL is able to match or sometimes outperform the performance of PPO while out-performing purely offline techniques like DPO (Rafailov et al., 2023).
Are there simpler algorithms that scale to modern RL applications?
@@
Our answer is REBEL: an algorithm that reduces the problem of reinforcement learning to solving a sequence of squared loss regression problems on iteratively collected datasets. The regression problems directly use policies to predict the difference in rewards. This allows us to eliminate the complexity of value functions, avoid heuristics like clipping, and scale easily to problems in both language modeling and image generation. Our key insight is that regressing relative rewards via policies directly on a sequence of iteratively collected datasets implicitly enables policy improvement.
@@
Rather than being a heuristic, REBEL comes with strong guarantees in theory and can be seen as a strict generalization of classical techniques (e.g., NPG) in reinforcement learning. Furthermore, REBEL cleanly incorporates offline datasets when available, can be extended to robustly handle intransitive preferences (Swamy et al., 2024), and empirically out-performs techniques like PPO and DPO (Rafailov et al., 2023) in language generation and has a faster convergence with a similar asymptotic performance in image generation. More explicitly, our key contributions are four-fold:
@@
    We connect REBEL to classical RL methods. We show that REBEL is a generalization of the foundational Natural Policy Gradient (NPG, Kakade (2001)) algorithm – applying the Gauss-Newton algorithm to the sequence of regression problems that REBEL solves recovers NPG. However, by instead applying simpler first-order optimization techniques, we are able to avoid computing the Fisher Information Matrix and enjoy a variance reduction effect. Thus, REBEL can be understood as a generalization of NPG while being much more scalable.
@@
    We analyze the convergence properties of REBEL. We prove via a direct reduction-based analysis that as long as we can solve the regression problem well at each iteration, we will be able to compete with any policy covered by the iteratively collected datasets (matching the strongest known results in the agnostic RL). These problems involve predicting the difference in rewards between trajectories in our dataset. We expect this problem to be well-solved in practice because our class of regressors is isomorphic to a class of policies that is highly expressive for the applications we consider (i.e. flexible Transformer models).
@@
    We evaluate REBEL both on language modeling and image generation tasks. We find that the on-policy version of REBEL outperforms PPO and DPO on language modeling and has similar performance for image generation tasks. On the TL;DR summarization task, we show REBEL scales well by finetuning a 6.9B parameter model. For text-guided image generation, REBEL optimizes a consistency model that converges to a similar performance as PPO.
In short, REBEL is a simple and scalable algorithm that enjoys strong theoretical guarantees and empirical performance. We believe it is a suitable answer to the question raised above.
@@ 2 REBEL: REgression to RElative REward Based RL
We first outline the notation used throughout the paper.
2.1 Notation
We consider the Contextual Bandit formulation (Langford and Zhang, 2007) of RL which has been used to formalize the generation process of models like LLMs (Rafailov et al., 2023; Ramamurthy et al., 2022; Chang et al., 2023) and Diffusion Models (Black et al., 2023; Fan et al., 2024; Oertell et al., 2024) due to the determinism of the transitions. More explicitly, in the deterministic transition setting, explicit states are not required as they can be equivalently represented by a sequence of actions. Furthermore, the entire sequence of actions can be considered as a single “arm” in a bandit problem with an exponentially large action space.
We denote by $(x,y)$ a prompt/response pair with $x\in\mathcal{X}$ as a prompt and $y\in\mathcal{Y}$ as a response (e.g., a sequence of tokens, or in general a sequence of actions). We assume access to a reward function $r(x,y)$ from which we can query for reward signals (the exact form of $r$ does not need to be known). Querying $r$ at $(x,y)$ will return a scalar $r(x,y)$ measuring the quality of the response. Such a reward function could be a pre-defined metric (e.g., Rouge score against human responses) or it could be learned from an offline human demonstration or preference data (e.g., the RLHF paradigm (Christiano et al., 2017; Ziegler et al., 2020)), as explored in our experiments.
Denote by $\pi\in\mathcal{X}\mapsto\Delta(Y)$, a policy (e.g. LLM) that maps from a prompt $x$ to a distribution over the response space $\mathcal{Y}$. We use $\rho$ to denote the distribution over prompts (i.e. initial states / contexts) $x$. Throughout the paper, we use $\pi_{\theta}(y|x)$ to denote a parameterized policy with parameter $\theta$ (e.g., a neural network policy). At times we interchangeably use $\pi_{t}$ and $\pi_{\theta_{t}}$ when it is clear from the context. We emphasize that while we focus on the bandit formulation for notation simplicity, the algorithms proposed here can be applied to any deterministic MDP where $x$ is the initial state and the trajectory $y$ consists of the sequence of actions.
@@ 2.1 Notation
$\pi_{t+1}=\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{x,y\sim\pi(\cdot|x)}r(x,y)-\frac{1}{\eta}\mathbb{E}_{x}\text{KL}\left(\pi(\cdot|x)||\pi_{t}(\cdot|x)\right).$ (1)
Intuitively, this can be thought of asking for the optimizer to fine-tune the policy $\pi_{t+1}$ according to $r$ while staying close to some baseline policy $\pi_{t}$.
@@ 2.2 Deriving REBEL: REgression to RElative REward Based RL
As first pointed out by Rafailov et al. (2023), observe that we can invert Eq. 2 and write the reward as a function of the policy, i.e. the “DPO Trick”:
@@
As soon as $\mathcal{X}$ and $\mathcal{Y}$ become large, we can no longer guarantee the above expression holds exactly at all $(x,y)$ and therefore need to turn our attention to choosing a policy such that Eq. 3 is approximately true. We propose using a simple square loss objective between the two sides of Eq. 3 to measure the goodness of a policy, i.e. reducing RL to a regression problem:
$\left(r(x,y)-\frac{1}{\eta}\left(\ln(Z(x))+\ln\left(\frac{\pi_{t+1}(y|x)}{\pi_{t}(y|x)}\right)\right)\right)^{2}.$ (4)
Unfortunately, this loss function includes the partition function $Z(x)$, which can be challenging to approximate over large input / output domains. However, observe that $Z(x)$ only depends on $x$ and not $y$. Thus, if we have access to paired samples, i.e. $(x,y)$ and $(x,y^{\prime})$, we can instead regress the difference in rewards to eliminate this term from our objective:
@@
Of course, we need to evaluate this loss function on some distribution of samples. In particular, we propose using an on-policy dataset $\mathcal{D}_{t}=\{x,y,y^{\prime}\}$ with $x\sim\rho,y\sim\pi_{t}(\cdot|x),y^{\prime}\sim\mu(\cdot|x)$, where $\mu$ is some base distribution. The base distribution $\mu$ can either be a fixed offline dataset (e.g. the instruction fine-tuning dataset) or $\pi_{t}$ itself. Thus, the choice of base distribution $\mu$ determines whether REBEL is hybrid or fully online. Putting it all together, we arrive at our core REBEL objective:
$\sum_{(x,y,y^{\prime})\in\mathcal{D}_{t}}\left((r(x,y)-r(x,y^{\prime}))-\frac{1}{\eta}\left(\ln\left(\frac{\pi_{t+1}(y|x)}{\pi_{t}(y|x)}\right)-\ln\left(\frac{\pi_{t+1}(y^{\prime}|x)}{\pi_{t}(y^{\prime}|x)}\right)\right)\right)^{2}.$ (6)
To recap, given a pair of completions $y,y^{\prime}$ to a prompt $x$, REBEL attempt to fit the relative reward
$r(x,y)-r(x,y^{\prime})$ (7)
by optimizing over a class of predictors of the form
$\frac{1}{\eta}\left(\ln\frac{\pi_{\theta}(y|x)}{\pi_{\theta_{t}}(y|x)}-\ln\frac{\pi_{\theta}(y^{\prime}|x)}{\pi_{\theta_{t}}(y^{\prime}|x)}\right).$ (8)
Critically, observe that if we were able to perfectly solve this regression problem, we would indeed recover the optimal solution to the KL-constrained RL problem we outlined in Eq. 1. While the above update might seem somewhat arbitrary at first glance, it has deep connections to prior work in the literature that illuminate its strengths over past techniques. We now discuss some of them.
Algorithm 1 REgression to RElative REward Based RL (REBEL)
1:  Input: Reward $r$, policy class $\Pi=\{\pi_{\theta}\}$, base distribution $\mu$, learning rate $\eta$
2:  Initialize policy $\pi_{\theta_{0}}$.
3:  for $t=0$ to $T-1$ do
4:     // Base distribution $\mu$ can either be an offline dataset or $\pi_{t}$.
5:     Collect dataset $\mathcal{D}_{t}=\{x,y,y^{\prime}\}$ where $x\sim\rho,y\sim\pi_{t}(\cdot|x),y^{\prime}\sim\mu(\cdot|x)$
6:     Solve square loss regression problem:
$\theta_{t+1}=\mathop{\mathrm{argmin}}_{\theta}\sum_{(x,y,y^{\prime})\in\mathcal{D}_{t}}\left(\frac{1}{\eta}\left(\ln\frac{\pi_{\theta}(y|x)}{\pi_{\theta_{t}}(y|x)}-\ln\frac{\pi_{\theta}(y^{\prime}|x)}{\pi_{\theta_{t}}(y^{\prime}|x)}\right)-\left(r(x,y)-r(x,y^{\prime})\right)\right)^{2}$ (9)
7:  end for

### v2

@@ Abstract
While originally developed for continuous control problems, Proximal Policy Optimization (PPO) has emerged as the work-horse of a variety of reinforcement learning (RL) applications, including the fine-tuning of generative models. Unfortunately, PPO requires multiple heuristics to enable stable convergence (e.g. value networks, clipping), and is notorious for its sensitivity to the precise implementation of these components. In response, we take a step back and ask what a minimalist RL algorithm for the era of generative models would look like. We propose REBEL, an algorithm that cleanly reduces the problem of policy optimization to regressing the relative reward between two completions to a prompt in terms of the policy, enabling strikingly lightweight implementation. In theory, we prove that fundamental RL algorithms like Natural Policy Gradient can be seen as variants of REBEL, which allows us to match the strongest known theoretical guarantees in terms of convergence and sample complexity in the RL literature. REBEL can also cleanly incorporate offline data and be extended to handle the intransitive preferences we frequently see in practice. Empirically, we find that REBEL provides a unified approach to language modeling and image generation with stronger or similar performance as PPO and DPO, all while being simpler to implement and more computationally efficient than PPO. When fine-tuning Llama-3-8B-Instruct, REBEL achieves strong performance in AlpacaEval 2.0, MT-Bench, and Open LLM Leaderboard.
@@ 1 Introduction
The generality of the reinforcement learning (RL) paradigm is striking: from continuous control problems (Kalashnikov et al., 2018) to, more recently, the fine-tuning of generative models (Stiennon et al., 2022; Ouyang et al., 2022), RL has enabled concrete progress across a variety of decision-making tasks. Specifically, when it comes to fine-tuning generative models, Proximal Policy Optimization (PPO, Schulman et al. (2017)) has emerged as the de-facto RL algorithm of choice, from language models (LLMs) (Ziegler et al., 2020; Stiennon et al., 2022; Ouyang et al., 2022; Touvron et al., 2023) to generative image models (Black et al., 2023; Fan et al., 2024; Oertell et al., 2024).
@@
If we take a step back however, it is odd that we are using an algorithm designed for optimizing two-layer networks for continuous control tasks from scratch, even though we are now fine-tuning generative models with billions of parameters. In the continuous control setting, the randomly initialized neural networks and the possible stochasticity in the dynamics necessitate variance reduction through a learned value function as a baseline (Schulman et al., 2015), while clipping updates is important to limit distribution shift from iteration to iteration (Kakade and Langford, 2002). This means that when applied to generative model fine-tuning, we need to store four models in memory simultaneously (the policy, the reference policy, the critic, and the reward model), each with billions of parameters. Furthermore, we often add a KL regularization to the base model for fine-tuning, making explicit clipping unnecessary nor advisable, as pointed out by Ahmadian et al. (2024). Even outside of the generative modeling context, PPO is notorious for the wide range of performances measured, with differences being attributed to seemingly inconsequential implementation details (Henderson et al., 2019; Engstrom et al., 2020). This begs the question: are there simpler algorithms that better scale to modern RL applications?
@@
Figure 1: We present REBEL: a simple and scalable RL algorithm that performs policy optimization via iteratively regressing the difference in rewards in terms of the policy, allowing us to eliminate much of the complexity (e.g. value functions, clipping) of algorithms like PPO (Schulman et al., 2017). We apply REBEL to problems in both image generation and language modeling and find that despite its conceptual and implementation-level simplicity, REBEL is able to match or sometimes outperform the performance of PPO while out-performing purely offline techniques like DPO (Rafailov et al., 2023). REBEL also achieves strong performance on common benchmarks such as AlpacaEval when fine-tuning a Llama-3-8B model.
@@
Our answer is REBEL: an algorithm that reduces the problem of RL to solving a sequence of squared loss regression problems on iteratively collected datasets. Each regression problem directly uses the policy to predict the difference in rewards. This allows us to eliminate the complexity of using value functions, avoids heuristics like clipping, and scales easily to problems in both language modeling and image generation. Our key insight is that a regressor that can predict the difference in rewards between trajectories in a dataset implicitly captures an improved policy.
@@
Rather than being a mere heuristic, REBEL comes with strong guarantees in theory and can be seen as a strict generalization of classical techniques (e.g., NPG) in reinforcement learning. Furthermore, REBEL cleanly incorporates offline datasets when available, can be extended to robustly handle intransitive preferences (Swamy et al., 2024), empirically out-performs techniques like PPO and DPO (Rafailov et al., 2023) in language generation, and has a faster convergence with a similar asymptotic performance in image generation. When fine-tuning a Llama-3-8B model, REBEL also demonstrates very competitive performance on the following benchmarks simultaneously: AlpacaEval 2.0 (length-controlled win-rate 30.1%), MT-bench (average 8.16), and Open LLM Leaderboard (average 68.2), without additional information such as online GPT4 queries. More explicitly, our key contributions are four-fold:
@@
    We connect REBEL to classical RL methods. We show that REBEL is a generalization of the foundational Natural Policy Gradient (NPG, Kakade (2001)) algorithm – applying the Gauss-Newton algorithm to the sequence of regression problems that REBEL solves recovers NPG. However, by instead applying simpler first-order optimization techniques, we are able to avoid computing the Fisher Information Matrix and enjoy a variance reduction effect under the setting of finite data. Thus, REBEL can be understood as a generalization of NPG while being much more scalable.
@@
    We analyze the convergence properties of REBEL. We prove via a direct reduction-based analysis that as long as we can solve the regression problem well at each iteration, we will be able to compete with any policy covered by the iteratively collected datasets (matching the strongest known results in the agnostic RL setting). These regression problems involve predicting the difference in rewards between trajectories in our dataset. We expect these regressions to be well-solved in practice because our class of regressors is isomorphic to a class of policies that is highly expressive for the applications we consider (i.e. flexible Transformer models).
@@
    We evaluate REBEL both on language modeling and image generation tasks. On the TL;DR summarization task, we show REBEL scales well by finetuning a 6.9B parameter model, and demonstrate that REBEL outperforms various baselines including PPO, DPO, REINFORCE, and RLOO. When fine-tuning a Llama-3-8B model, REBEL demonstrates very competitive performance on AlpacaEval, MT-bench, and Open LLM Leaderboard. For text-guided image generation, REBEL optimizes a consistency model that converges faster to a similar performance faster than PPO. Implementation of REBEL can be found at https://github.com/ZhaolinGao/REBEL, and models trained by REBEL can be found at https://huggingface.co/Cornell-AGI.
We begin by formalizing the preference fine-tuning setup before deriving our core algorithmic technique.
@@ 2 REBEL: REgression to RElative REward Based RL
We consider the contextual bandit formulation (Langford and Zhang, 2007) of RL which has been used to formalize the generation process of models like LLMs (Rafailov et al., 2023; Ramamurthy et al., 2022; Chang et al., 2023) and Diffusion Models (Black et al., 2023; Fan et al., 2024; Oertell et al., 2024) due to the determinism of the transitions. More explicitly, in the deterministic transition setting, explicit states are not required as they are isomorphic to the sequence of actions. Furthermore, the entire sequence of actions can be considered as a single “arm” in a bandit problem with an action space that scales exponentially in size with the horizon of the problem.
We denote by $(x,y)$ a (prompt, response) pair, where $x\in\mathcal{X}$ is the prompt and $y\in\mathcal{Y}$ is the response (e.g. a sequence of tokens, or in general a sequence of actions). We assume access to a reward function $r(x,y)$ from which we can query for reward signals (the exact form of $r$ does not need to be known). Querying $r$ at $(x,y)$ will return a scalar $r(x,y)$, measuring the quality of the prompt completion. Such a reward function could be a pre-defined metric (e.g., Rouge score against human responses) or a learned model from an offline human demonstration or preference data (e.g. the RLHF paradigm (Christiano et al., 2017; Ziegler et al., 2020)), as we focus on in our experiments.
Denote by $\pi\in\mathcal{X}\mapsto\Delta(Y)$ a policy (e.g. an LLM) that maps from a prompt $x$ to a distribution over the response space $\mathcal{Y}$. We use $\rho$ to denote the distribution over prompts (i.e. initial states / contexts) $x$ and $\pi_{\theta}(y|x)$ to denote a policy with parameter $\theta$. At times, we interchangeably use $\pi_{t}$ and $\pi_{\theta_{t}}$ when it is clear from the context. We emphasize that while we focus on the bandit formulation for notational simplicity, the algorithms proposed here can be applied to any deterministic MDP where $x$ is the initial state and the trajectory $y$ consists of the sequence of actions.
Algorithm 1 REgression to RElative REward Based RL (REBEL)
1:  Input: Reward $r$, policy class $\Pi=\{\pi_{\theta}:\theta\in\Theta\}$, base distribution $\mu$, learning rate $\eta$
2:  Initialize policy $\pi_{\theta_{0}}$.
3:  for $t=0$ to $T-1$ do
4:     // Base distribution $\mu$ can either be an offline dataset or $\pi_{t}$.
5:     Collect dataset $\mathcal{D}_{t}=\{x,y,y^{\prime}\}$ where $x\sim\rho,y\sim\pi_{t}(\cdot|x),y^{\prime}\sim\mu(\cdot|x)$.
6:     Solve square loss regression problem:
$\theta_{t+1}=\mathop{\mathrm{argmin}}_{\theta\in\Theta}\sum_{(x,y,y^{\prime})\in\mathcal{D}_{t}}\left(\frac{1}{\eta}\left(\ln\frac{\pi_{\theta}(y|x)}{\pi_{\theta_{t}}(y|x)}-\ln\frac{\pi_{\theta}(y^{\prime}|x)}{\pi_{\theta_{t}}(y^{\prime}|x)}\right)-\left(r(x,y)-r(x,y^{\prime})\right)\right)^{2}.$ (1)
7:  end for
@@
$\pi_{t+1}=\mathop{\mathrm{argmax}}_{\pi\in\Pi}\mathbb{E}_{x,y\sim\pi(\cdot|x)}r(x,y)-\frac{1}{\eta}\mathbb{E}_{x}\text{KL}\left(\pi(\cdot|x)||\pi_{t}(\cdot|x)\right).$ (2)
Intuitively, this can be thought of asking for the optimizer to fine-tune the policy $\pi_{t+1}$ according to $r$ while staying close in terms of action distribution to some baseline policy $\pi_{t}$.
@@ 2.1 Deriving REBEL: REgression to RElative REward Based RL
As observed by Rafailov et al. (2023), we can invert Eq. 3 and write the reward in terms of the policy:
@@
As soon as $\mathcal{X}$ and $\mathcal{Y}$ become large, we can no longer guarantee the above expression holds exactly at all $(x,y)$ and therefore need to turn our attention to choosing a policy such that Eq. 4 is approximately true. We propose using a simple square loss objective between the two sides of Eq. 4 to measure the goodness of a policy, i.e. reducing RL to a regression problem: $\left(r(x,y)-\frac{1}{\eta}\left(\ln(Z(x))+\ln\left(\frac{\pi_{t+1}(y|x)}{\pi_{t}(y|x)}\right)\right)\right)^{2}$. Unfortunately, this loss function includes the partition function $Z(x)$, which can be challenging to approximate over large input / output domains. However, observe that $Z(x)$ only depends on $x$ and not $y$. Thus, if we have access to paired samples, i.e. $(x,y)$ and $(x,y^{\prime})$, we can instead regress the difference in rewards to eliminate this term:
@@
Of course, we need to evaluate this loss function on some distribution of samples. In particular, we propose using an on-policy dataset $\mathcal{D}_{t}=\{x,y,y^{\prime}\}$ with $x\sim\rho,y\sim\pi_{t}(\cdot|x),y^{\prime}\sim\mu(\cdot|x)$, where $\mu$ is some base distribution. The base distribution $\mu$ can either be a fixed offline dataset (e.g. the instruction fine-tuning dataset) or $\pi_{t}$ itself. Thus, the choice of base distribution $\mu$ determines whether REBEL is hybrid or fully online. Putting it all together, we arrive at our core REBEL objective in Eq. 1. Critically, observe that if we were able to perfectly solve this regression problem, we would indeed recover the optimal solution to the KL-constrained RL problem we outlined in Eq. 2.
2.2 Extending REBEL to Stochastic MDPs
In this section, we demonstrate that REBEL can be applied to any stochastic MDP as long as we have the ability to reset to the initial states. Let $x$ denote the initial state $s_{0}$ of the MDP and $y$ represent a sequence of states and actions $(a_{0},s_{1},a_{1},...,a_{H-1},s_{H})$ where $H$ is the sequence length and $a_{h}\sim\pi_{\theta}(\cdot|s_{h})$. The probability of trajectory $y$ given $x$ is $\pi_{\theta}(y|x)=\prod_{h}\pi_{\theta}(a_{h}|s_{h})P(s_{h+1}|s_{h},a_{h})$ where $P$ is the transition probability. As in a policy gradient, observe that the transition probability cancels out since REBEL always uses probability ratios: $\pi_{\theta}(y|x)/\pi_{\theta_{t}}(y|x)=\prod_{h}\pi_{\theta}(a_{h}|s_{h})/\pi_{\theta_{t}}(a_{h}|s_{h})$. By resetting to the initial state $s_{0}$, we can generate another trajectory $y^{\prime}$ from $\pi_{\theta}$ with length $H$, leading to the following objective:
$\theta_{t+1}=\mathop{\mathrm{argmin}}_{\theta}\sum_{(x,y,y^{\prime})\in\mathcal{D}_{t}}\left(\frac{1}{\eta}\left(\sum_{h}\ln\frac{\pi_{\theta}(a_{h}|s_{h})}{\pi_{\theta_{t}}(a_{h}|s_{h})}-\sum_{h^{\prime}}\ln\frac{\pi_{\theta}(a^{\prime}_{h^{\prime}}|s^{\prime}_{h^{\prime}})}{\pi_{\theta_{t}}(a^{\prime}_{h^{\prime}}|s^{\prime}_{h^{\prime}})}\right)-\left(r(x,y)-r(x,y^{\prime})\right)\right)^{2}$
where $s_{0}=s^{\prime}_{0}=x$. We emphasize again that we only require resetting to the initial states. Note that similar to REINFORCE, the trajectory-level rewards could be high-variance, perhaps introducing the need for critic-based variance reduction (Konda and Tsitsiklis, 1999; Mnih et al., 2016; Haarnoja et al., 2018). We leave the experimental validation of REBEL in stochastic MDPs to future work.